# Optimizing Cycle Life Prediction of Lithium-ion Batteries via a Physics-Informed Model

### v1

@@ Abstract
Keywords— electric vehicles, state of health, Arrhenius law, discharge voltage curve, self-attention, multi-stage training
@@ 1 Introduction
Previous literature on predicting cycle lifetimes of batteries is rich. Data-driven models [25, 4, 2, 21], which focus on using machine learning techniques to identify trends in how batteries degrade, have been thoroughly studied, with models ranging from linear models [12] to neural networks [4, 14, 15] to support vector regression [26]. These types of models are agnostic to the mechanisms of degradation, but they can be difficult to fine-tune. Physics-based models, which rely upon knowledge in how a battery degrades over time, have also been well-studied. These methods usually rely upon cell chemistry [19, 10, 24] or analyzing how the material in the electrodes changes over time [5, 9]. However, since batteries can be charged/discharged in a variety of environments, this hinders how descriptive physical models by themselves can be. Recently, hybrid models combining physics knowledge with a data-driven approach have been proposed [22, 6, 8, 11] to combine the advantages of both approaches.
@@ 2 Data Processing
We utilize the public dataset provided by Severson et al. [12]. This dataset comprises 124 lithium-ion phosphate/graphite battery cells, each with a nominal capacity of 1.1 ampere-hours (Ah) [1]. The cells are cycled (repeatedly charged/discharged) until end of life, which is defined as the point where the effective capacity of the battery has dropped to 80% of its nominal value. The ratio of the effective and nominal capacity is commonly referred to as the state of health. All battery cells in this dataset are cycled under fast-charging conditions in a constant-temperature environment; however, the charging policy dictating the specific charge rate schedule differs from cell to cell.
In our work, we preserve the train/test/secondary test data split in Severson et al. [12], allowing for a direct comparison between our result and theirs. The primary test set was obtained using the same batch of cells as the train set and similar charge policies, therefore we use it to evaluate the model’s ability to interpolate in the input space. On the other hand, the secondary test set was obtained from a different batch of cells and using significantly different scheduling, and we use it to examine the model’s ability to extrapolate. Ensuring our model can generalize to different batteries makes for an advantage over prior work [11, 14] that uses a random split of cells into train/validation/test sets.
For each cell, three forms of data are recorded:
@@
Figure 1: (a) Discharged capacity, $Q_{d}(t)$, and voltage, $V(t)$, measured for one battery over the course of a charge-discharge cycle. The discharge portion of the cycle is boxed in black. (b) The discharge-voltage curve, $Q_{d}(V)$, for the same battery. N.B.: the x-axis in (b) is flipped to reflect voltage decreasing over the course of the discharging process.
@@
Evolution of $Q_{d}(V)$ over cycles can be exploited to predict battery lifetime. Severson et al. [12] observe that more dramatic early-cycle curve sagging occurs for batteries with low lifetimes than for batteries with high lifetimes, as visualized in Figure 2(a, b). To capture the phenomenon of curve sagging, Severson et al. propose taking the discharge-voltage curves for cycles 100 and 10 and computing the difference between the two. This new curve, calculated as $Q_{d,100}(V)-Q_{d,10}(V)$, is denoted $\Delta Q_{100-10}(V)$. Figure 2(c,d) shows that $\Delta Q_{100-10}(V)$ succinctly capture the difference in curve sagging behavior between two batteries of different cycle lives, and Figure 2(e) shows a clear linkage between curve shape and cycle life across all batteries in the dataset. In particular, batteries with shorter cycle lives exhibit more ample dips in the $\Delta Q_{100-10}(V)$ curve.
@@ 3.1 Physics-Based Model
It is well known that as a lithium-ion battery is cycled, other chemical processes occur in the battery that affects long term cyclability. Notably, the SEI (solid electrode interphase) is formed on the surface of the anode within the first five charging cycles, impeding electron movement in the battery. Although poorly understood, it is believed that SEI formation has an impact on capacity loss [17]. Hence it is important to consider the impact of SEI formation when studying capacity loss. In response, we construct a model of the capacity curves in the Severson et al. dataset via an Arrhenius Law, which is commonly associated with chemical processes akin to SEI formation [18, 13, 3].
We now introduce an equation describing the process of capacity loss. Wang et al. (2011) [18] have shown that for different charge rates, the true capacity loss $Q_{\text{loss}}$ can be approximated as
$\hat{Q}(x)_{\text{loss}}=B\exp\left(-\frac{E_{a}}{RT}\right)x^{z}$ (2)
@@
for constants $B$ and $z$, where $\hat{Q}_{\text{loss}}$ represents the percentage of capacity loss, $E_{a}$ is the activation energy, $R$ the gas constant, $T$ the absolute temperature, and $x$ the cycle number. Note that the temperature highly resembles an Arrhenius Law, which is commonly associated with a thermally activated chemical process, such as SEI layer formation during cycling [18, 13, 3]. Equation (2) accounts for SEI formation that produces undesired side effects in batteries. In addition, there is also a power law with respect to the cycle number, consistent with previous findings of the rate of lithium consumption at the negative electrode [18, 13, 20, 3]. In short, this equation accounts for SEI formation and is consistent with the resulting mechanism of active lithium consumption in the presence of the SEI layer.
@@
We adapt Equation (2) in three ways. Firstly, we noticed that typical values of the constant $B$ are found to have a very large order of magnitude, which introduces numerical instability and overflow issues in further calculations. We mitigate this by predicting $\ln B$ instead of $B$ itself. Secondly, the average temperatures reached during testing did not vary greatly across cycles. Treating $T$ as invariant allows us to incorporate the original exponential factor from Equation (2) into the previously described constant. This has the subtle advantage that $E_{a}$ need not be known anymore.
We lastly adopt the addition of a constant $C$ by shifting up the predicted curves such that their first point matches the first point of the ground truth capacity loss curve. Specifying such constant reduces one free variable and makes fitting easier. Thus, we may reduce Equation (2) to one of the form
$\hat{Q}(x)_{\text{loss}}=e^{A}x^{B}+Q(0).$ (3)
@@
We then fit Severson et al’s capacity loss curves with equations of the form in Equation (3) via least squares regression. Figure 3(a-c) illustrates three capacity loss curves with substantially different cycle lives and their best fit curves. Not only are the curves themselves a good fit, but the cycle lives as predicted by our best fit curves are remarkably similar to the actual cycle lives. Cycle life $\ell$ is the point where $\hat{Q}_{\text{loss}}(\ell)=0.2$, calculated from the best fit curve as
$\ell=[e^{-A}(0.2-Q(0))]^{1/B},$ (4)
@@
where 0.2 is used as the threshold capacity loss indicating end of life. Figure 3d plots true cycle life against predicted cycle life for all batteries in the dataset and demonstrates high goodness of fit, with $R^{2}=0.994$ and a RMSE of 28.6.
@@
As seen in Figure 3(d), our equation very accurately models the cycle lives of batteries, and henceforth we assume that the ground truth of capacity loss follows Equation (3), given values of $A$ and $B$. In other words, we assume that our predicted $\hat{Q}_{\text{loss}}\approx Q_{\text{loss}}$. Hence the second half of our hybrid model involves training a self-attention model to predict $\hat{A}$ and $\hat{B}$ from cycle input data of the first 100 cycles. Call this model $f$. Then given a vector $\mathbf{x}$ of early-cycle statistical quantities, our output variables are $(\hat{A},\hat{B})=f(\mathbf{x})$ and our predicted cycle life is ^$\hat{\ell}=[e^{-\hat{A}}(0.2-C)]^{1/\hat{B}}$. This model is illustrated in Figure 4.
Figure 4: Schematic of the physics-based model. One half utilizes an Arrhenius Law-inspired model to capture capacity loss curves. The other half utilizes self-attention to predict Arrhenius Law parameters from early-cycle data.
@@ 3.2 Self-Attention for Regression
We endeavor to predict the parameters of the capacity loss curve based on early-cycle data, when the full capacity loss curve is not known. When limited to early-cycle capacity loss data, we cannot fit and extrapolate the full curve with good fidelity using least squares. Instead, we employ a self-attention mechanism to learn complex, nonlinear relations between the capacity loss curve and other measurements available during a battery’s early operation. As presented in [7], there exists an equivalence between the self-attention mechanism and a support vector regression (SVR) problem formulation. This implies that any problem where an SVR model performs well can potentially be improved by employing an attention-based architecture.
Let the input sequence of $N$ features derived from early-cycle operation data for the $i$-th battery be denoted $\mathbf{x_{i}}:=[x_{i}^{(1)},\dots,x_{i}^{(N)}]^{T}\in\mathbb{R}^{N\times 1}$. We compute the standard query matrix Q, key matrix K, and value matrix V from the self-attention mechanism via the following transformations:
$\mathbf{Q}_{i}=\mathbf{x}_{i}\mathbf{W}^{T}_{Q}$
$\mathbf{K}_{i}=\mathbf{x}_{i}\mathbf{W}^{T}_{K}$
$\mathbf{V}_{i}=\mathbf{x}_{i}\mathbf{W}^{T}_{V},$
@@ 3.3 Feature Selection
Finally, we select the best features for prediction. In Supplementary Table 1, we list all the features we obtain from the charge and discharge curves as explained in Section 2. Our main task is to carefully select the features that are most closely related to our target variables, $\hat{A}$ and $\hat{B}$. To accomplish this, we analyze the correlation between each feature and the target variables. We prioritize features with strong positive or negative correlations, as they are more likely to provide accurate predictions.
We use Spearman’s correlation coefficient to ascertain which features are most correlated with $\hat{A}$ and $\hat{B}$. Spearman’s correlation coefficient is a statistical measure that assesses the strength and direction of the monotonic relationship between two variables. It is particularly useful when the relationship between variables is nonlinear. In our analysis, we select the top five features with the highest correlation coefficients. The results of our correlation analysis using this approach are visually represented in Figure 5. Note that these features are DeltaQ_logVars, DeltaQ_logMin, DeltaQ_logMean, slope_capacity_fade_2_100, and slope_capacity_91_100. These five features were used to train the self-attention model as explained in Section 3.2.
Figure 5: Feature correlation of the selected features with equation parameters.
@@ 3.4 Model Training
where $w_{A}$ and $w_{B}$ are tunable hyperparameters. The parameter loss function is smoother and leads to fewer numerical issues than Equation (6), producing stable results when training from a random initialization. In contrast, we observe that training only with the cycle life loss function leads to exploding gradients and inconsistent behavior.
However, parameter loss is not always indicative of the accuracy of cycle life predictions. Due to the high nonlinearity of the capacity loss curve, it is possible for two sets of parameter estimates to incur equal parameter loss but produce dramatically different cycle life predictions (Supplementary Figure 1). Consequently, we follow up coarse training with a fine-tuning stage under low learning rate that trains on cycle life loss.
@@
The choice of optimal hyperparameters are based upon the existing literature, and such parameters can be consulted in the Github repository associated with our work, available upon request.
@@ 4 Results
We initiate the model training phase with a basic regularized linear framework. The baseline model is similar to the model that performed the best in Severson et al.’s findings and utilizes sklearn’s ElasticNet model [12]. For hyperparameter tuning, we vary alpha and l1_ratio on a logscale over $10^{0}$ and $10^{1}$ and $10^{-5}$ and $10^{2}$, respectively. We find that the best model had a primary test RMSE of 398.98 cycles and a secondary test RMSE of 455.4 cycles.
The results of our model evaluation are presented in Figure 6. We compare the performance of the elastic net baseline with that of the self-attention model in terms of train RMSE, primary test RMSE, and secondary test RMSE. Results from the self-attention model are divided into two stages, Stage 01 and Stage 02, corresponding to the two phases of training.
Note that the errors achieved by self-attention after two stages of training are significantly better than the elastic net baseline; further, they are comparable with Severson et al.’s original errors [12]. After two training stages, we were able to improve primary test RMSE from 398.87 to 127.83 cycles and secondary test RMSE from 455.4 to 179.92 cycles.
Figure 7 illustrates the true versus predicted cycle lives (a) as well as true and predicted capacity curves for sample batteries (b-d) using the fully trained self-attention model. We notice that the predicted capacity loss curves do indeed fit the behavior of the ground truth curves across train, primary test, and secondary test batteries.
Further, recall that Severson et al.’s full model utilizing multiple features achieves a primary test RMSE of 118 and a secondary test RMSE of 214. Our primary test error is quite similar, but our secondary test error improves upon theirs by 30 cycles (over a 15% improvement). Given that our model captures more information on how a battery degrades over time, we conclude that our model serves as an appealing alternative to predict battery cycle life, especially in cases where we wish to have the flexibility to define cycle life differently based on state of health.

### v2

@@ 1 Introduction
Previous literature on predicting cycle lifetimes of batteries is rich. Data-driven models [27, 4, 2, 23], which focus on using machine learning techniques to identify trends in how batteries degrade, have been thoroughly studied, with models ranging from linear models [13] to neural networks [4, 16, 17] to support vector regression [28]. These types of models are agnostic to the mechanisms of degradation, but require significant effort to fine-tune, as their performance depends heavily on hyperparameter optimization, feature selection, and the quality of the training data. Physics-based models, which rely upon knowledge in how a battery degrades over time, have also been well-studied. These methods usually rely upon cell chemistry [22, 10, 26] or analyzing how the material in the electrodes changes over time [5, 9]. However, since batteries can be charged/discharged in a variety of environments, this hinders how descriptive physical models by themselves can be. Recently, hybrid models combining physics knowledge with a data-driven approach have been proposed [24, 6, 8, 11] to combine the advantages of both approaches.
@@
Using the hybrid model, we are able to predict battery capacity loss curves. Cycle life is defined as the point where the effective capacity of the battery has dropped to a percentage of its nominal capacity. In the literature, 80% is typically the threshold, but other studies have considered 85% [12, 14]. With entire capacity loss curves, we are able to predict cycle life for all other thresholds without retraining, while models that are specifically trained with the 80% threshold must be retrained. This accounts for conditions that may require a different threshold for cycle life, such as the make of the battery and the presence of non-ideal operating conditions.
@@ 2 Data Processing
We utilize the public dataset provided by Severson et al. [13]. This dataset comprises 124 lithium-ion phosphate/graphite battery cells, each with a nominal capacity of 1.1 ampere-hours (Ah) [1]. The cells are cycled (repeatedly charged/discharged) until end of life, which we take as the point where the effective capacity of the battery has dropped to 80% of its nominal value. The ratio of the effective and nominal capacity is commonly referred to as the state of health. All battery cells in this dataset are cycled under fast-charging conditions in a constant-temperature environment; however, the charging policy dictating the specific charge rate schedule differs from cell to cell.
In our work, we preserve the train/test/secondary test data split in Severson et al. [13], allowing for a direct comparison between our result and theirs. The primary test set was obtained using the same batch of cells as the train set and similar charge policies; therefore we use it to evaluate the model’s ability to interpolate in the input space. On the other hand, the secondary test set was obtained from a different batch of cells and using significantly different scheduling, and we use it to examine the model’s ability to extrapolate. Ensuring our model can generalize to different batteries makes for an advantage over prior work [11, 16] that uses a random split of cells into train/validation/test sets.
For each cell, four forms of data are recorded:
@@
Figure 1: (a) Discharged capacity, $Q_{d}(t)$, and voltage, $V(t)$, measured for one battery over the course of a charge-discharge cycle. The discharge portion of the cycle is boxed in black. (b) The discharge-voltage curve, $Q_{d}(V)$, for the same battery. Note that the x-axis in (b) is flipped to reflect voltage decreasing over the course of the discharging process.
@@
The evolution of $Q_{d}(V)$ over cycles can be exploited to predict battery lifetime. Severson et al. [13] observe that more dramatic early-cycle curve sagging occurs for batteries with low lifetimes than for batteries with high lifetimes, as visualized in Figure 2(a, b). To capture the phenomenon of curve sagging, Severson et al. [13] propose taking the discharge-voltage curves for cycles 100 and 10 and computing the difference between the two. This new curve, calculated as $Q_{d,100}(V)-Q_{d,10}(V)$, is denoted $\Delta Q_{100-10}(V)$. Figure 2(c,d) shows that $\Delta Q_{100-10}(V)$ succinctly capture the difference in curve sagging behavior between two batteries of different cycle lives, and Figure 2(e) shows a clear linkage between curve shape and cycle life across all batteries in the dataset. In particular, batteries with shorter cycle lives exhibit more ample dips in the $\Delta Q_{100-10}(V)$ curve.
@@ 3.1 Physics-Based Model
It is well known that as a lithium-ion battery is cycled, other chemical processes occur in the battery that affects long term cyclability. Notably, the SEI (solid electrode interphase) is formed on the surface of the anode within the first five charging cycles, impeding electron movement in the battery. Although poorly understood, it is believed that SEI formation has an impact on capacity loss [19]. Hence it is important to consider the impact of SEI formation when studying capacity loss. In response, we construct a model of the capacity curves in the Severson et al. [13] dataset via an Arrhenius Law, which is commonly associated with chemical processes akin to SEI formation [20, 15, 3].
We now introduce an equation describing the process of capacity loss. Wang et al. [20] have shown that for different charge rates, the true capacity loss $Q_{\text{loss}}$ can be approximated as [^1] Two adaptations have been made to the original Equation (2) from [20]. Firstly, we replaced the dependency on time with a dependency on cycle number. This is acceptable since all cycles last equal amounts of time. Secondly, we changed the notation of the parameters to not conflict with notations we introduce next.
$\hat{Q}_{\text{loss}}(x)=D\exp\left(-\frac{E_{a}}{RT}\right)x^{B}$ (2)
@@
for constants $B$ and $D$, where $\hat{Q}_{\text{loss}}$ represents the percentage of capacity loss, $E_{a}$ is the activation energy, $R$ the gas constant, $T$ the absolute temperature, and $x$ the cycle number. Note that the temperature dependence in the exponential term resembles the Arrhenius Law, which is commonly associated with a thermally activated chemical process, such as SEI layer formation during cycling [20, 15, 3]. Equation (2) aims to account for SEI formation that produces undesired side effects in batteries. In addition, there is also a power law with respect to the cycle number, consistent with previous findings of the rate of lithium consumption at the negative electrode [20, 15, 21, 3]. In short, this equation seeks to describe SEI formation and is consistent with the resulting mechanism of active lithium consumption in the presence of the SEI layer.
We adapt Equation (2) in three ways, while maintaining its physical interpretation. In our dataset, we observe that cells typically show an initial capacity loss even at cycle 0, which can be attributed to various factors, including calendar aging during storage and formation cycles. To account for this, we introduce a constant term $C$ that represents this initial capacity loss. We use the nominal capacity as our reference than the measured capacity at cycle 0, as this provides a consistent baseline across all cells and allows for more robust comparison between different cells, particularly when initial conditions may vary due to storage time or formation cycles.
@@
Next, we observe that the average temperatures reached during testing did not vary greatly across cycles. Treating $T$ as invariant allows us to incorporate the original exponential factor from Equation (2) into the constant $D$. This has the advantage that $E_{a}$ need not be known anymore.
Lastly, we note that typical values of the constant $D$ have a very large order of magnitude, which introduces numerical instability and overflow issues in further calculations. We mitigate this by predicting $A\equiv\ln D$ instead of $D$ itself.
Thus, we may reduce Equation (2) to one of the form given by Equation (3). By defining $C$ to be equal to the initial capacity loss $Q_{\text{loss}}(0)$, we arrive at a functional form with just two parameters, $A$ and $B$.
$\hat{Q}_{\text{loss}}(x)=e^{A}x^{B}+C$ (3)
@@
We then fit the capacity loss curves in [13] to Equation (3) via least squares regression. Figure 3(a-c) illustrates three capacity loss curves with substantially different cycle lives and their best fit curves. Not only are the curves themselves a good fit, but the cycle lives as predicted by our best fit curves are remarkably similar to the actual cycle lives. Cycle life $\ell$ is the point where $\hat{Q}_{\text{loss}}(\ell)=0.2$, calculated from the best fit curve as
$\ell=[e^{-A}(0.2-C)]^{1/B},$ (4)
@@
where 0.2 is used as the threshold capacity loss indicating end of life. Figure 3(d) plots true cycle life against predicted cycle life for all batteries in the dataset and demonstrates high goodness of fit, with $R^{2}=0.994$ and a RMSE of 28.6 cycles.
@@
As seen in Figure 3(d), our equation very accurately models the cycle lives of batteries, and henceforth we assume that the ground truth of capacity loss follows Equation (3), given values of $A$ and $B$. In other words, we assume that $\hat{Q}_{\text{loss}}\approx Q_{\text{loss}}$.
Now the other half of our hybrid model involves training a self-attention model to predict $\hat{A}$ and $\hat{B}$ from cycle input data of the first 100 cycles, which will be explained in detail in the section below. Call this model $f$. Then given a vector $\mathbf{x}$ of early-cycle statistical quantities, our output variables are $(\hat{A},\hat{B})=f(\mathbf{x})$ and our predicted cycle life is ^$\hat{\ell}=[e^{-\hat{A}}(0.2-C)]^{1/\hat{B}}$. This model is illustrated in Figure 4.
Figure 4: Schematic of the physics-based model. One half utilizes an Arrhenius Law-inspired model to capture capacity loss curves. The other half utilizes a self-attention layer to predict Arrhenius Law parameters from early-cycle data.
@@ 3.2 Self-Attention for Regression
We endeavor to predict the parameters of the capacity loss curve based on early-cycle data, when the full capacity loss curve is not known. When limited to early-cycle capacity loss data, we cannot fit and extrapolate the full curve with good fidelity using least squares. Instead, we employ a single self-attention block to learn complex, nonlinear relations between the capacity loss curve and other measurements available during a battery’s early operation. As presented in Nguyen et al. [7], there exists an equivalence between the self-attention mechanism and a support vector regression (SVR) problem formulation. This implies that any problem where an SVR model performs well can potentially be improved by employing an attention-based architecture.
Let the input sequence of $N$ features derived from early-cycle operation data for the $i$-th battery be denoted $\mathbf{z_{i}}:=[z_{i}^{(1)},\dots,z_{i}^{(N)}]^{T}\in\mathbb{R}^{N\times 1}$, where each $z(j)_{i}$ represents a measurement feature such as voltage, current, and temperature . We compute the standard query matrix Q, key matrix K, and value matrix V from the self-attention mechanism via the following transformations:
$\mathbf{Q}_{i}=\mathbf{z}_{i}\mathbf{W}^{T}_{Q}$
$\mathbf{K}_{i}=\mathbf{z}_{i}\mathbf{W}^{T}_{K}$
$\mathbf{V}_{i}=\mathbf{z}_{i}\mathbf{W}^{T}_{V},$
@@ 3.3 Feature Selection
Finally, we select the best features for prediction from the available cycle summary features and discharge curve characteristics described in Section 2. From the cycle summary data, we extract capacity fade characteristics by calculating slopes over different cycle windows. From the discharge-voltage curves $Q_{d}(V)$, we derive statistical features including the variance, minimum, and mean of the $\Delta$Q curves, following the approach introduced by Severson et al. (2019). Additionally, we calculate logarithmic transformations of these quantities to better capture the nonlinear relationships in the data.
Our main task is to carefully select the features that are most closely related to our target variables, $\hat{A}$ and $\hat{B}$. To accomplish this, we analyze the correlation between each feature and the target variables. We prioritize features with strong positive or negative correlations, as they are more likely to provide accurate predictions. We use Spearman’s correlation coefficient to ascertain which features are most correlated with $\hat{A}$ and $\hat{B}$. In our analysis, we select the top five features with the highest correlation coefficients. The results of our correlation analysis using this approach are visually represented in Figure 5 and the full correlation coefficients can be found in Appendix A.
Note that these features are:
  1. 1.
    DeltaQ_logVars: The log of the variance of the $\Delta Q_{100-10}(V)$ curve.
  2. 2.
    DeltaQ_logMin: The log of the min of the $\Delta Q_{100-10}(V)$ curve.
  3. 3.
    DeltaQ_logMean: The log of the mean of the $\Delta Q_{100-10}(V)$ curve.
  4. 4.
    slope_capacity_fade_2_100: Capacity fade rate from cycle 2 to 100. This is the rate that the capacity fade curve decreases from cycle 2 to cycle 100.
  5. 5.
    slope_capacity_91_100: Capacity fade rate from cycles 91 to 100. This is the rate that the capacity fade curve decreases from cycle 91 to cycle 100.
These five features were used to train the self-attention model as explained in Section 3.2.
Figure 5: Feature correlation scores of the five features most correlated with $\hat{A}$ and $\hat{B}$.
@@ 3.4 Model Training
where $w_{A}$ and $w_{B}$ are tunable hyperparameters. The parameter loss function is smoother and leads to fewer numerical issues than Equation (6), producing stable results when training from a random initialization. In contrast, we observe that training only with the original cycle life loss function leads to exploding gradients and inconsistent behavior.
However, parameter loss is not always indicative of the accuracy of cycle life predictions. Due to the high nonlinearity of the capacity loss curve, it is possible for two sets of parameter estimates to incur equal parameter loss but produce dramatically different cycle life predictions. Consequently, we follow up coarse training with a fine-tuning stage under low learning rate that trains on cycle life loss.
@@
Training took place with $800$ epochs and a learning rate of $10^{-3}$ in step 1 and with $3000$ epochs and a learning rate of $5\times 10^{-5}$ in step 2. The choice of optimal hyperparameters are based upon the existing literature, and such parameters can be consulted in the Github repository associated with our work, available upon request. A full comparison of the relevant training curves can be found in Appendix B and Appendix C.
@@ 4 Results
We initiate the model training phase with a baseline model using a regularized linear framework. For the baseline model, we take sklearn’s ElasticNet model that performed the best in Severson et al. [13]. However, to ensure comparable results, we train the baseline model on the features chosen in Section 3.3. We further change the output of the baseline model to predict $\hat{A}$ and $\hat{B}$ in order reconstruct the entire capacity curve. The results of the baseline model are shown in Figure 6.
For hyperparameter tuning, we vary alpha and l1_ratio on a logscale over $10^{0}$ and $10^{1}$ and $10^{-5}$ and $10^{2}$, respectively. We find that the best model had a primary test RMSE of 398.98 cycles and a secondary test RMSE of 455.4 cycles. We see that in this case, the linear assumptions of elastic net limit its ability to capture nonlinear degradation patterns and parameterize capacity loss curves.
Next, we trained the hybrid model (also referred to as the self-attention model) to predict the parameters $\hat{A}$ and $\hat{B}$. The results of our model evaluation are presented in Figure 6, which compares the performance of the elastic net baseline with that of the self-attention model in terms of the train RMSE, primary test RMSE, and secondary test RMSE. Results from the self-attention model are divided into two stages, Stage 1 and Stage 2, corresponding to the two phases of training.
Note that the errors achieved by self-attention after two stages of training are significantly better than the elastic net baseline; further, they are comparable with the original RMSE errors in Severson et al. [13]. After two training stages, we were able to improve the primary test RMSE from 398.87 to 127.83 cycles and the secondary test RMSE from 455.4 to 179.92 cycles.
Figure 7 illustrates the true versus predicted cycle lives (a) as well as true and predicted capacity curves for sample batteries (b-d) using the fully trained self-attention model. We notice that the predicted capacity loss curves do indeed fit the behavior of the ground truth curves across train, primary test, and secondary test batteries. Further, recall that the full model in Severson et al. [13] utilizing multiple features achieves a primary test RMSE of 118 and a secondary test RMSE of 214. Our primary test error is quite similar, but our secondary test error improves upon theirs by 30 cycles (over a 15% improvement). Given that our model captures more information on how a battery degrades over time, we conclude that our model serves as an appealing alternative to predict battery cycle life, especially in cases where one wishes to have the flexibility to consider a different definition of cycle life.